Given the original demand and supply equations for period 1, Qd = 120 - 6P and Qs = 12 + 4P solve the following:


Problem 1: The new and original price equilibrium given that the new demand and supply equations are Qd = 130 - 6P and Qs = 20 + 4P


Problem 2: The new and original quantity equilibrium given that the new demand and supply equations are Qd = 130 - 6P and Qs = 20 + 4P


Problem 3: The change in price equilibrium given that the new demand and supply equations are Qd = 130 - 6P and Qs = 20 + 4P


Problem 4: The change in quantity equilibrium given that the new demand and supply equations are Qd = 130 - 6P and Qs = 20 + 4P



1
Expert's answer
2022-03-08T12:55:34-0500

1.The original demand equation = "Q"d "=120-6P"

The original supply equation ="Q"s= "12+4P"

At equilibrium "Q"d"=Q"s

"120-6P=12+4P"

"120-12=4P+6P"

"=108=10P"

"=108\/10=10P\/10"

"=10.8=P"

"\therefore" "P=10.8"

The original price is 10.8.

The new demand equation = "Q"d "=130-6P"

The new supply equation ="Q"s"=20+4P"

At equilibrium "Q"d"=Q"s

"130-6P=20+4P"

"=130-20=4P+6P"

"=110=10P"

"=110\/10=10P\/10"

"=11=P"

"\therefore" "P=11"

The new price is 11.

2.The original quantity demanded="Q"d"=120-6P"

"=120-6*10.8"

"=55.2"

The original quantity supplied "=Q"s"=12+4P"

"=" "12+4*10.8"

"=" "55.2"

The new quantity demanded= "Q"d"=130-6P"

"=130-6*11"

"=64"

The new quantity supplied = 64

3.The change in price equilibrium = "New price -Original price"

"= 11-10.8"

"=0.2"

4.The change in quantity equilibrium "=New quantity - Original quantity"

"=64-55.2"

"=8.8"